OpenStudy (anonymous):

hi, i need help, i dont understand this problem, how would i go about doing this? Which of the following is not an equivalent form of the compound inequality: x + 18 <= 10 and x + 18 >= 5 * A number line with a closed circle on -13, shading to the left, and a closed circle on -8, shading to the right. *10 >= x + 18 >= 5 * -13 <= x <= -8 * A number line with a closed circle on -13, a closed circle on -8, and shading in between. please tell me the answer and exactly how you got the answer so i can learn.

OpenStudy (anonymous):

third one i guess

OpenStudy (accessdenied):

You should justify those answers. Simple answers do not necessarily tell how to get them in Math unfortunately.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

well thanks for the answer, though it would've been better to explain how you got it, for my benefit. ty anyways though

OpenStudy (anonymous):

closing it now.

OpenStudy (accessdenied):

Well, the first thing to realize is that "x + 18 <= 10 and x + 18 >= 5" can be rewritten as "5 <= x + 18 <= 10." This is logically equivalent to (B.) if we reverse the order (10 >= x + 18 >= 5). So, (B.) is NOT the answer! If we solve for the variable interval by subtracting 18 in both cases: x + 18 <= 10, x <= -8; x + 18 >= 5, x >= -13. Once again, if we write this in the compact form, we get "-13 <= x <= -8," which is answer choice (C.), so (C.) is also not the answer. We lastly observe how we plot this. The closed circles go on -13 and -8. We then shade the space between them, which is what (D.) says; so (D.) is not the answer. This is to the right of -13 and to the left of -8. It appears that (A.) is actually incorrect! We cannot shade to the right of -13 because this is the values less than -13, like -16, etc.